# Does Diffusion Beat GAN in Image Super Resolution?

### v1

@@ Abstract
There is a prevalent opinion in the recent literature [32] that Diffusion-based models outperform GAN-based counterparts on the Image Super Resolution (ISR) problem. However, in most studies, Diffusion-based ISR models were trained longer and utilized larger networks than the GAN baselines. This raises the question of whether the superiority of Diffusion models is due to the Diffusion paradigm being better suited for the ISR task or if it is a consequence of the increased scale and computational resources used in contemporary studies. In our work, we compare Diffusion-based and GAN-based Super Resolution under controlled settings, where both approaches are matched in terms of architecture, model and dataset size, and computational budget. We show that a GAN-based model can achieve results comparable to a Diffusion-based model. Additionally, we explore the impact of design choices such as text conditioning and augmentation on the performance of ISR models, showcasing their effect on several downstream tasks. We will release the inference code  [^1] https://github.com/yandex-research/gan_vs_diff_sr and weights of our scaled GAN.
@@ 1 Introduction
In contrast, DDMs, which have gained traction more recently, provide a robust framework for image generation, reliably covering multi-modal data distributions. However, the Diffusion ability to yield high-quality and diverse outputs comes with a cost of high computational load during training and inference.
At the moment, the literature suggests that Diffusion models become a dominant source of state-of-the-art methods in generative modeling, including ISR ([41, 28, 55, 53]). However, a comprehensive comparison between GAN and Diffusion models in the context of ISR under controlled conditions still needs to be explored. In particular, GAN-based methods that modern state-of-the-art methods are compared with are typically represented by Real-ESRGAN [48], which has orders of magnitude fewer parameters, training dataset, and computational budget.
This work aims to thoroughly compare the GAN and DDM frameworks for Image Super Resolution. Given the unique strengths of both GANs and Diffusion models, understanding their comparative performance in ISR tasks is crucial. This comparison requires a rigorous and controlled experimental setup to ensure fairness, i.e., standardized variables such as the size of the training dataset, model complexity, and controlled allocation of computational resources.
Furthermore, we explore the influence of various design choices on the performance and generalization of these models. In particular, we investigate the value of practically useful Real-ESRGAN [48] augmentations and validate the necessity of incorporation of textual conditioning (e.g., image captions). To sum up, our main contributions are:
  • We thoroughly compare GAN and DDM for Image Super Resolution and reveal that scaled appropriately, Diffusion and GAN result in similar Super Resolution quality in various downstream tasks.
  • We investigate the influence of textual conditioning in the form of non-SR-specific image captions and find that it has no significant effect on both frameworks.
@@ 3 Methodology
Since our goal is to compare GAN and Diffusion Super Resolution under a comparable setup, we use the same architecture for both paradigms considered. Specifically, we adopt a network similar to $256\times 256\rightarrow 1024\times 1024$ SR model from Imagen ([40]). It acts in the pixel domain and is based on Efficient U-Net architecture [35]. This model has $600-700$M trainable parameters, depending on the configuration. Details about architecture are provided in Appendix B.
@@
$L_{\mathrm{diff}}=\mathbb{E}_{I^{HR}\sim p(I^{HR}),I^{LR}\sim p(I^{LR}),\epsilon\in\mathcal{N}(0,1),t\in\mathcal{U}[0,1]}\|\epsilon-\epsilon_{\theta}(z_{t},I^{LR},t)\|_{2}^{2},z_{t}=\alpha_{t}I^{HR}+\sigma_{t}\epsilon$ (4)
@@
GAN-based upscalers produce high-resolution images in a single forward pass, whereas Diffusion models iteratively denoise input noise conditioned on a low-resolution image. For that, we adopt DPM-Solver++ [29] as the one providing an excellent trade-off between generation quality and efficiency. We set the order of the solver to be two and chose the multistep method for sampling. All upscaled images are generated with 13 sampling steps. Our evaluations showed that a further increase in the number of sampling steps does not lead to visually noticeable image quality gains.
@@ 4.1 Experimental setup
We train our models on a large-scale proprietary dataset of 17 million text-image pairs. The dataset consists of $1024\times 1024$ px images with exceptional image quality, high image-text relevance, and English captions. The high dataset quality is achieved by a rigorous filtering process described in detail in Appendix E.
@@
We follow prior work in the choice of hyperparameters [48, 43] and observe that the training with these choices is stable and converges fast enough. We fix the batch size for GAN and Diffusion to ensure fairness of comparison on the number of samples seen during training and adjust the learning rate accordingly. Further increasing the learning rate with the goal of speeding up training leads to training instability and poor final results. For both models, the batch size is set to 512 for training on crops and 128 for full-resolution images. We adopt a constant linear rate with a linear warmup phase. The models are trained with Adam  [16] optimizer without weight decay. More details are provided in Appendix C.
While most of the related works adopt a predefined number of training iterations, we set the training duration dynamically and terminate the training only when the visual quality of the images upscaled by adjacent model checkpoints becomes indistinguishable from the humans’ perspective. Since the changes become smaller throughout the training, we use an exponentially spaced grid with checkpoints from the following list of thousands of training iterations: [10, 20, 40, 60, 100, 140, 220, $\ldots$], where we increase the interval between adjacent two checkpoints by a factor of two. Then, we estimate the human annotator preference using the Side-by-Side (SbS) evaluation for the current and previous training steps. If there is no improvement for three subsequent measurements, we conclude that the model performance is saturated and proceed to the next stage. We note that the training loss and full-reference metrics may continue to improve, but the improvement is imperceptible to human observers.
We emphasize that by design, the aim of SbS comparisons is to track the model progress rather than assess the final performance. We observed that SbS between two relatively high-quality models can be very noisy due to the ambiguity and high level of subjectivity in assessing small differences in upscaling quality. For example, if one model tends to imperfectly sharpen the region and the second model to smooth that, it may not be obvious which one is closer to the ground truth. Because of that, the victory of one model over another may flip from evaluation to evaluation. On the other hand, we observed that models obtained from different checkpoints of the same training procedure act more similarly i. e., models that tend to over-smooth images in the early stages of training do less of it later. Therefore, we only adopt SbS comparisons to determine the optimal training duration for both SR paradigms, leaving the judgment of the final quality for the Reader and automated metrics. The SbS comparison setup is described and explained in greater detail in Appendix D
@@
In addition, we investigate the robustness of ISR models on out-of-distribution synthetic data in Appendix F. For that, we provide a qualitative comparison between GAN and Diffusion models together with the baselines on the Super Resolution of downscaled images produced by the YaART [17] and SDXL [33] models using prompts from DrawBench [40] and YaBasket [17] prompt sets.
@@ 4.2 Training dynamics
Evaluation results are presented in Figure 1. We treat models as equal if $p$-value on SbS comparison $>0.05$ following common practice. More details are in Appendix D.
@@ 4.3 Performance comparison
To assess the performance of SR models, we report classic full-reference metrics PSNR, SSIM [46], LPIPS [57] and recent no-reference CLIP-IQA [47] metric in Table 1. Figure 2 presents several representative visual samples. As mentioned above, we do not employ the SbS comparison for performance estimates due to high subjectivity, resulting in noisy results when both models are close.
@@ Results
GAN-based and Diffusion-based SR models yield upsampled images of approximately the same quality with respect to evaluation metrics. Both approaches produce visually aesthetic samples with sharp edges, small details, and apparent textures. Our SR models perform comparable to or better than the baselines from the literature.
@@ 4.4 Impact of text conditioning
We train text-text-conditioned modelsmodels with the same training setup used for unconditional models. Our experiments cover two text encoders of different nature: our internal CLIP-like model called XL and UMT5 Text-conditioned models have slightly more parameters than unconditional models due to additional transformer blocks. We conduct an SbS comparison between text-conditioned and unconditional models for both paradigms at all stages of training. According to the results in Figure 3 human annotators show preference neither for text-conditional models nor for unconditional.
@@ 4.5 ISR for a complex degradation model and the impact of training augmentations
We observe that $L_{1}$ pretraining and training with the adversarial loss for GAN models converge with the same number of training steps as for training without augmentations (see Figure 4). However, with Real-ESRGAN augmentations, the Diffusion SR model requires even more extended training, and its performance saturates only after 940k iterations. The augmentations appear to noticeably slow down the convergence of the Diffusion model. We also observe that the problem of color shift that is known for under-trained Diffusion SR models [20] becomes more apparent in augmentation-heavy pipelines.
@@ Results
We show a couple of samples for the dataset with Real-ESRGAN-augmented pipeline in Figure 5. Diffusion and GAN-based SR yield upscaled images of similar quality.
@@ 5 Ablations
Table 4: Human preferences for GAN SR models without finetuning on full resultion images and with finetuning.
@@ 6 Limitations
Another exciting question left for future work is the study of scaling behavior for both GAN and Diffusion SR paradigms concerning the model size and amount of training data and compute.
@@ 7 Conclusion
Appendix A Qualitative Comparison

### v2

@@ Abstract
There is a prevalent opinion  [^1] See, for example, [33]. that diffusion-based models outperform GAN-based counterparts in the Image Super Resolution (ISR) problem. However, in most studies, diffusion-based ISR models employ larger networks and are trained longer than the GAN baselines. This raises the question of whether the high performance stems from the superiority of the diffusion paradigm or if it is a consequence of the increased scale and the greater computational resources of the contemporary studies. In our work, we thoroughly compare diffusion-based and GAN-based Super Resolution models under controlled settings, with both approaches having matched architecture, model and dataset sizes, and computational budget. We show that a GAN-based model can achieve results comparable or superior to a diffusion-based model. Additionally, we explore the impact of popular design choices, such as text conditioning and augmentation on the performance of ISR models, showcasing their effect in several downstream tasks. We will release the inference code  [^2] https://github.com/yandex-research/gan_vs_diff_sr and weights of our scaled GAN.
@@ 1 Introduction
In contrast, more contemporary DDMs provide a robust framework for image generation, reliably covering multi-modal data distributions. Yet, the diffusion ability to yield high-quality and diverse outputs comes with a cost of high computational load during training and inference.
At the moment, the literature suggests that diffusion models have become a dominant source of state-of-the-art methods in generative modeling, including ISR ([43, 29, 59, 57]). However, a comprehensive comparison between GAN and diffusion models in the context of ISR under controlled conditions still needs to be explored. In particular, GAN-based methods that modern state-of-the-art methods are compared with are typically represented by Real-ESRGAN [52], which has orders of magnitude fewer parameters, training data samples, and computational budget.
This work aims to thoroughly compare GAN and DDM frameworks for Image Super Resolution. Given the unique strengths of both GANs and diffusion models, understanding their comparative performance in ISR tasks is crucial. This comparison requires a rigorous and controlled experimental setup to ensure fairness, i.e., standardized variables such as the size of the training dataset, model complexity, and controlled allocation of computational resources.
Furthermore, we explore the influence of various design choices on the performance and generalization of these models. In particular, we investigate the value of practically useful Real-ESRGAN [52] augmentations and validate the necessity of incorporating textual conditioning (e.g., image captions). The contributions of our work are the following:
  • We thoroughly compare GAN and DDM in Image Super Resolution problem: if scaled appropriately, GAN-based models surpass quality of diffusion-based ones in classical SR and yield similar quality in SR in the Wild.
  • We investigate the influence of textual conditioning in the form of non-ISR-specific image captions and find that it has no significant effect on both frameworks.
@@ 3 Methodology
Since our goal is to compare GAN and diffusion Super Resolution under a comparable setup, we use the same architecture for both paradigms considered. Specifically, we adopt a network similar to $256\times 256\rightarrow 1024\times 1024$ SR model from Imagen ([42]). It acts in the pixel domain and is based on Efficient U-Net architecture [37]. This model has $600-700$M trainable parameters, depending on the configuration. Details about architecture are provided in Appendix C.
@@
$L_{\mathrm{diff}}=\mathbb{E}_{I^{HR}\sim p(I^{HR}),I^{LR}\sim p(I^{LR}),\epsilon\sim\mathcal{N}(0,1),t\sim\mathcal{U}[0,1]}\|\epsilon-\epsilon_{\theta}(z_{t},I^{LR},t)\|_{2}^{2},z_{t}=\alpha_{t}I^{HR}+\sigma_{t}\epsilon$ (4)
@@
GAN-based upscalers produce high-resolution images in a single forward pass, whereas diffusion models iteratively denoise input noise conditioned on a low-resolution image. For that, we adopt DPM-Solver++ [30] as the one providing an excellent trade-off between generation quality and efficiency. We set the order of the solver to be two and chose the multistep method for sampling. All upscaled images are generated with 13 sampling steps. Our evaluations showed that a further increase in the number of sampling steps does not lead to visually noticeable image quality gains H.1.
@@ 4.1 Experimental setup
We train our models on a large-scale proprietary dataset of 17 million text-image pairs. The dataset consists of $1024\times 1024$ px images with exceptional image quality, high image-text relevance, and English captions. The high dataset quality is achieved by a rigorous filtering process described in detail in Appendix  F.
@@
We follow prior work in the choice of hyperparameters [52, 46] and observe that the training with these choices is stable and converges fast enough. We fix the batch size for GAN and diffusion to ensure fairness of comparison on the number of samples seen during training and adjust the learning rate accordingly. Further increasing the learning rate with the goal of speeding up training leads to training instability and poor final results. For both models, the batch size is set to 512 for training on crops and 128 for full-resolution images. We adopt a constant linear rate with a linear warmup phase. The models are trained with Adam  [15] optimizer without weight decay. More details are provided in Appendix D.
While most of the related works adopt a predefined number of training iterations, we set the training duration dynamically and terminate the training only when the visual quality of the images upscaled by adjacent model checkpoints becomes indistinguishable from the humans’ perspective. Since the changes become smaller throughout the training, we use an exponentially spaced grid with checkpoints from the following list of thousands of training iterations: [10, 20, 40, 60, 100, 140, 220, $\ldots$], where we increase the interval between adjacent two checkpoints by a factor of two. Then, we estimate the human annotator preference using the Side-by-Side (SbS) evaluation for the current and previous training steps. If there is no improvement for three subsequent measurements, we conclude that the model performance is saturated and proceed to the next stage. We note that the training loss and full-reference metrics may continue to improve, but the improvement is imperceptible to human observers. In our work, user preferences are used both for early stopping and model evaluation. The SbS comparison setup is described and explained in greater detail in Appendix E.
@@
In addition, we investigate the robustness of ISR models on out-of-distribution synthetic data in Appendix G. For that, we provide a qualitative comparison between GAN and diffusion models together with the baselines on the Super Resolution of downscaled images produced by the YaART [16] and SDXL [34] models using prompts from DrawBench [42] and YaBasket [16] prompt sets.
@@ 4.2 Training Dynamics
Evaluation results are presented in Figure 1. We treat models as equal if $p$-value on SbS comparison $>0.05$ following common practice. More details are in Appendix  E.
@@ 4.3 Performance comparison
To assess the performance of SR models, we report classic full-reference metrics PSNR, SSIM [50], LPIPS [62] and recent no-reference CLIP-IQA [51] metric in Table 1. Figure 2 presents several representative visual samples. Additionally, we conduct side-by-side (SbS) comparisons between different models and present user preference results in Table 2 and Table 5. We calculate the number of wins of one model over another plus the number of draws divided by two, which we denote as the win rate. The win rates are color-coded: Green if the first model is statistically better than the second, red if the first model is statistically worse, and black otherwise. Details about the statistical significance criteria are provided in Appendix E.
@@ Results
GAN-based SR model is equal or better than diffusion with respect to the majority of evaluation metrics. Yet both approaches produce visually aesthetic samples with sharp edges, small details, and apparent textures. Our SR models are competitive to the baselines from the literature. According to the user preferences (Table 2) GAN-based model outperforms any of the other models considered. We note that full-reference metrics do not always agree with each other and SbS comparison results.
@@
Table 2: SbS comparison between GAN-based and diffusion-based SR with each other and current state-of-the-art. The values in the table are win rates of Model 1 over Model 2. Green corresponds to statistical advantage, red to statistical disadvantage, black to statistical indifference between two models.
Model 1 Model 2 Diff (ours) GAN (ours) SUPIR RealESRGAN DiffBIR ResShift
Diff (ours) x 9.0 37.7 85.2 51.0 46.3
GAN (ours) 91.0 x 62.7 94.6 69.0 86.0
@@ 4.4 Impact of Text Conditioning
We train text-conditioned emodels with the same training setup used for unconditional models. Our experiments cover two text encoders of different nature: our internal CLIP-like model called XL and UMT5 Text-conditioned models have slightly more parameters than unconditional models due to additional transformer blocks. We conduct an SbS comparison between text-conditioned and unconditional models for both paradigms at all stages of training. According to the results in Figure 3 human annotators show preference neither for text-conditional models nor for unconditional.
@@ 4.5 ISR for a Complex Degradation Model and the Impact of Training Augmentations
We observe that $L_{1}$ pretraining and training with the adversarial loss for GAN models converge with the same number of training steps as for training without augmentations (see Figure  4). However, with Real-ESRGAN augmentations, the diffusion SR model requires even more extended training, and its performance saturates only after 940k iterations. The augmentations appear to noticeably slow down the convergence of the diffusion model.
@@ Results
We show a couple of samples for the dataset with Real-ESRGAN-augmented pipeline in Figure 5. GAN-based and diffusion-based SR models yield samples of similar quality. GAN-based SR from our study appears to be the best in terms of full-reference metrics. At the same time, human annotators show preference for SR models with diffusion priors (DiffBIR, SUPIR). We speculate that presence of rich semantic information is more beneficial in the presence of degradations.
@@
Table 5: SbS comparison between GAN-based and diffusion-based SR with each other and current state-of-the-art. The values in the table are win rates of Model 1 over Model 2. Green corresponds to statistical advantage, red to statistical disadvantage, black to statistical indifference between two models.
Model 1 Model 2 Diff (ours) GAN (ours) SUPIR RealESRGAN DiffBIR ResShift
Diff (ours) x 34.0 18.4 71.7 29.3 72.1
GAN (ours) 66.0 x 30.0 66.2 30.6 80.2
@@
4.6 Artifacts
Both SR paradigms suffer from specific artifacts. GAN-based SR models are prone to oversharpening and generating unnatural textures. On the other hand, diffusion-based SR is usually inferior at generating high-frequency details. In addition, diffusion-based SR models tend to generate pale, dim images when compared to the original. We provide visualizations of these common artifacts in Figure 13 and Figure 14.
@@ 5 Ablations
Table 6: Human preferences for GAN SR models without finetuning on full resolution images and with finetuning.
@@
6 Efficiency
Another important aspect from a practical perspective is the efficiency of the Super Resolution pipeline. While one approach may offer better Super Resolution quality compared to another, this advantage could be outweighed by a much larger inference cost and memory overhead. As one can observe from Table 8, models differ by orders of magnitude in terms of the total number of parameters involved in all stages of the inference workflow. Some approaches can be easily deployed on mobile devices, while others necessitate a compute accelerator with a large enough amount of memory.
We carried out inference measurements on an NVIDIA A100 GPU for upscaling a single image from $256\rightarrow 1024$ (batch size is 1) and report the latency in Table 8. GAN-based methods produce a high-resolution image given a low-resolution input in a fraction of a second, whereas diffusion-based SR with generative priors requires dozens of seconds.
Table 8: Runtime and number of parameters (during inference) of compared models. Some of the models are consisted of several stages with parameters distributed among them. The benchmarking and parameter-counting protocol are explained in greater detail in Appendix H.5.
Diff (ours) GAN (ours) SUPIR RealESRGAN DiffBIR ResShift
Runtime, (s) $3.20\pm 0.21$ $0.24\pm 0.02$ $18.11\pm 0.85$ $0.065\pm 0.003$ $24.69\pm 0.03$ $1.19\pm 0.03$
# Parameters, (M) 631 614 17846 17 1683 174
@@ 7 Limitations
Another interesting question not covered in our investigation is the study of scaling behavior for both GAN and diffusion SR paradigms with respect to the model size and amount of training data and compute. In preliminary experiments we observed that the training parameters adopted in our main training setup do not transfer immediately to other model sizes and have to be tuned for specific experimental setup.
@@ 8 Conclusion
Appendix A Background on generative paradigms
In this section, we provide a mathematical background for the generative models used in our study.
A.1 Generative Adversarial Networks
To produce samples that resemble a target distribution, one can train an auxiliary critic network, known as a discriminator, to distinguish generated data from real data. The discriminator, denoted as $D_{\theta}$, is optimized alongside the generator, denoted as $G_{\theta}$, to solve the following adversarial min-max problem:
$\min_{\theta_{G}}\max_{\theta_{D}}\mathbb{E}_{I^{out}\sim p(I^{out})}[\log D_{\theta_{D}}(I^{out})]+\mathbb{E}_{I^{in}\sim p(I^{in})}[\log(1-D_{\theta_{D}}(G_{\theta_{G}}(I^{in})))]$ (5)
where $I^{in}$ and $I^{out}$ are the input and output spaces respectively. For instance, $I^{in}$ can be the space of random noise or low-resolution images (for the Super Resolution problem) and $I^{out}$ may correspond to the space of natural images. In the case of successful training, the generator learns to produce samples that are difficult to classify by $D_{\theta_{D}}$ and resemble the target distribution.
This approach was introduced in [8] and subsequently became a leading generative paradigm for an extended period. GANs have demonstrated the capability to generate high-quality, photorealistic samples [14, 17, 18]. However, they are known to suffer from issues such as unstable training and mode collapse [49].
A.2 Diffusion Models
Diffusion models ([9]) have emerged as a promising technique in the landscape of generative modeling. These models are primarily inspired by concepts from nonequilibrium thermodynamics, leveraging the idea of diffusion processes to gradually transform simple distributions into complex data distributions and vice versa.
The forward process in diffusion models is a mechanism that systematically transforms structured data into noise through a series of discrete steps. Starting with an original data point $z_{0}$, this process involves a Markov chain where each step $z_{t}$ given $z_{t-1}$ is defined by a Gaussian distribution
$q(z_{t}|z_{t-1})=\mathcal{N}(z_{t};\sqrt{\alpha_{t}}z_{t-1},(1-\alpha_{t})I),$ (6)
with $\alpha_{t}$ controlling the noise level. As the process progresses, the data is incrementally corrupted, ensuring the distribution of $z_{t}$ given $z_{0}$ remains Gaussian,
$q(z_{t}|z_{0})=\mathcal{N}(z_{t};\sqrt{\gamma_{t}}z_{0},(1-\gamma_{t})I),$ (7)
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$. This results in $z_{t}$ becoming increasingly dominated by noise, converging to a standard Gaussian distribution as $t$ approaches $T$. The formulation allows for the derivation of the posterior $q(z_{t-1}|z_{t},z_{0})=\mathcal{N}(z_{t-1};\mu_{t},\sigma_{t}^{2}I)$, providing a basis for the reverse process that aims to denoise the data back to its original form.
The backward diffusion process, or reverse process, in diffusion models, is essential for generating structured data from noise by reversing the forward diffusion steps. Starting from a noise sample $z_{T}\sim\mathcal{N}(0,I)$, this process iteratively denoises the sample through a sequence of steps $z_{T},z_{T-1},\ldots,z_{0}$, transforming it back into structured data $z_{0}$. Each reverse step $p_{\theta}(z_{t-1}|z_{t})$ is modeled as a Gaussian distribution:
$p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t),\sigma_{\theta}(z_{t},t)I),$ (8)
where $\mu_{\theta}$ and $\sigma_{\theta}$ are parameterized by a neural network, trained to approximate the reverse of the forward noise addition. The training objective typically maximizes a variational lower bound on the data log-likelihood, encouraging the learned reverse process to closely match the true reverse process. The posterior distribution from the forward process $q(z_{t-1}|z_{t},z_{0})=\mathcal{N}(z_{t-1};\mu_{t},\sigma_{t}^{2}I)$ guides the reverse process’s learning, ensuring that each step effectively reduces the noise incrementally added during the forward process.
This objective can be reduced ([9]) to a noise prediction one. Given the original data point $z_{0}$ and its noisy version $z_{t}$ at step $t$, produced by the forward process $q(z_{t}|z_{0})=\mathcal{N}(z_{t};\sqrt{\gamma_{t}}z_{0},(1-\gamma_{t})I)$, the model is tasked with predicting the noise $\epsilon$ that was added. The neural network, parameterized by $\theta$, outputs the predicted noise $\epsilon_{\theta}(z_{t},t)$, aiming to match the actual noise added during the diffusion step. The training objective minimizes the mean squared error between the predicted noise and the true noise:
$L_{\text{diff}}=\mathbb{E}_{z_{0},\epsilon\sim\mathcal{N}(0,I),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t)\|^{2}\right],$ (9)
where $z_{t}=\sqrt{\gamma_{t}}z_{0}+\sqrt{1-\gamma_{t}}\epsilon$ and $\epsilon\sim\mathcal{N}(0,I)$.
Finally, the diffusion framework can be extended to conditional modeling via appropriate conditioning mechanisms. By incorporating conditional information $c$, such as class labels, text prompts or images, the model is guided to generate data that is contextually relevant and aligned with the specified conditions.
Appendix B Qualitative Comparison